OpenStudy (anonymous):

S(t)=68-20log(t+1), t≥0 What was the average score after 14 months? After what time t was the average score 40%?

hero (hero):

1. S(14) = 68 - 20log(14 + 1) 2. 40 = 68 - 20log(t + 1)

OpenStudy (anonymous):

Thank you so much for your help!

hero (hero):

You're welcome.